### v2

# LOCUS – LOcalization with Channel Uncertainty and Sporadic Energy

Abstract.
Accurate sound source localization (SSL) requires consistent multichannel data for reliable degree of arrival (DoA) estimation. However, intermittently powered batteryless systems often suffer from incomplete sensor data due to the stochastic nature of energy harvesting. Existing methods struggle with missing channels, leading to significant performance degradation. In this paper, we propose LOCUS, a novel deep learning-based system designed to recover corrupted features for SSL in batteryless systems. LOCUS addresses missing data by leveraging information entropy estimation and conditional interpolation, combining three modules: (1) Information-Weighted Focus (InFo), which identifies and quantifies corrupted data elements, (2) Latent Feature Synthesizer (LaFS), which synthesizes missing features, and (3) Guided Replacement (GRep), which intelligently replaces missing elements while preserving valid data. We demonstrate significant performance improvements using two datasets: DCASE and LargeSet, where LOCUS achieves up to 36.91% lower DoA error compared to existing methods. Real-world evaluations across three environments with intermittent power sources show a 25.87-59.46% improvement in performance when channels are stochastically missing. Additionally, we release a 50-hour multichannel dataset to support further research in SSL.
1. Introduction
Accurate degree of arrival (DoA) estimation relies on extracting temporal and spatial differences from multichannel data captured by multiple microphones. Both signal processing (Schmidt, 1986; DiBiase, 2000; Roy and Kailath, 1989) and deep learning (Xiao et al., 2015; Yalta et al., 2017; Chakrabarty and Habets, 2017) approaches use this multichannel data to localize sound sources effectively. In real-world scenarios, where complex multipath profiles and room impulse responses arise from signal reflections and refractions, deep learning methods excel by leveraging the richness of multichannel data to capture intricate patterns that traditional signal processing techniques may miss.
However, intermittently powered batteryless systems face significant challenges in maintaining full-channel data integrity due to the stochastic nature of energy harvesting. These systems, reliant on sporadic power availability, often struggle to consistently capture data from all channels, which is crucial for precise DoA estimation. Intermittent systems typically can only process short-lived events, lasting only a few hundred milliseconds, and are primarily limited to tasks such as classification (Majid et al., 2020a; Islam and Nirjon, 2020; Monjur et al., 2023). This constraint arises from the core design of intermittent systems, which rely on a fixed-size capacitor to store sufficient harvested energy for operation. During the system’s active phase, it consumes this stored energy to perform designated tasks, after which it enters a sleep phase when the energy is depleted. By selecting different capacitor sizes, the duration of the active and sleep phases can be adjusted. However, intermittent systems inevitably encounter one or more sleep phases when insufficient energy is harvested, resulting in incomplete or missing sensor data (Monjur et al., 2023). This lack of sensor data significantly degrades the system’s performance, particularly for longer events that last on the order of tens of seconds. As a result, these systems are not typically used for sound source localization (SSL), where complete multichannel data is essential.
Despite these challenges, batteryless systems hold great potential in applications such as environmental monitoring and wildlife tracking, where low-power, autonomous operation is essential. Even though these systems experience frequent power interruptions, they can still provide valuable data during operational periods.
Working with features instead of raw time-domain data is particularly important in batteryless systems. Extracted features like generalized cross-correlation (GCC) and spectral power encapsulate the most relevant information for sound source localization, reducing noise and data dimensionality. This not only improves computational efficiency but also minimizes energy consumption, which is critical in resource-constrained systems. Moreover, feature extraction can be implemented using low-power analog circuits in embedded systems, further reducing the energy footprint.
Existing solutions for addressing missing or corrupted data in sound source localization typically rely on statistical imputation techniques such as mean imputation (Allison, 2001), hot deck imputation (Little and Rubin, 2019), and multiple imputations (Rubin, 1996). While these methods can fill data gaps, they have notable limitations. Mean imputation oversimplifies by averaging values, hot deck imputation can introduce inaccuracies if data is poorly matched, and multiple imputations are computationally intensive and struggle with long missing data sequences (Hasan et al., 2021). Recent approaches using mutual information and conditional entropy (Vergara and Estévez, 2014) help identify corrupted data but fall short in recovering the information necessary for DNN-based sound source localization. Additionally, recent work on imputation-based data recovery integrated into DNN pipelines for batteryless systems (Monjur et al., 2023) is designed for single-channel input, making it inadequate for the multi-channel sound source localization needed in more complex systems.
The gap in current sound source localization systems lies in the inability to effectively recover missing channel data caused by power interruptions in intermittently powered systems. Without advanced methods to intelligently infer or reconstruct this information, localization accuracy suffers significantly. Frequent data loss, especially in batteryless systems relying on stochastic energy harvesting, presents a major challenge for precise localization.
This paper bridges the gap between faulty multi-channel microphone arrays and deep learning-based batteryless sound source localization (SSL) by proposing LOCUS. LOCUS combines information entropy estimation and conditional interpolation to identify corrupted portions in the feature set and estimate the true feature values. The system consists of three novel modules: (1) information-weighted focus (InFo), which estimates the information entropy for each element of the calculated features to understand the impact of corrupted data; (2) latent feature synthesizer (LaFS), which estimates missing information by leveraging the relationships derived from information entropy; and (3) guided replacement (GRep), which uses the information entropy from InFo to replace corrupted elements with interpolated features from LaFS, leaving unaffected elements intact. These reconstructed features are then fed into the DNN for sound source localization.
The main contributions of this work are as follows:
  • We propose LOCUS, a three-step information recovery system that leverages information entropy estimation and conditional interpolation to recover corrupted features, mitigating performance loss in deep learning-based sound source localization due to missing or faulty input channels.
  • We provide a mathematical formulation of the problem and establish a connection with information theory principles.
  • Beyond evaluation with multiple datasets, we conduct real-world evaluations in three different environments, where missing channels occur due to power intermittence from energy harvesting systems (solar and radio frequency). This demonstrates the practical applicability and robustness of LOCUS in real-world scenarios.
  • We release a new 6-channel acoustic dataset (LargeSet) comprising 50 hours of audio recorded in 10 different environments. This dataset supports further research and development in sound source localization.
Our evaluation shows significant performance gains, with LOCUS achieving 21.27% and 17.14% less average DoA error than $SoundSeive$(Monjur et al., 2023) on the DCASE (Politis et al., 2021) and LargeSet datasets, respectively. In real-world evaluations, LOCUS improves performance by 25.87-59.46% in scenarios with stochastically missing information. Additionally, LOCUS reduces the 90th percentile localization error by 25.52% in wireless source localization tasks using the WiFi channel state information (CSI) dataset (Ayyalasomayajula et al., 2020).
2. Managing Missing Data in Batteryless Microphone Arrays
Quantifying Missing Data (MDP). In a multi-channel microphone array, data from a sound source can be corrupted by missing or intermittent data from one or more channels, while others may remain unaffected. To standardize the measurement of missing data, we define the Missing Data Percentage (MDP), denoted as $p\%$. The MDP quantifies the proportion of the total audio that is corrupted. The MDP is calculated using the following formula:
$MDP=\frac{t_{1}+t_{2}+\cdots+t_{n}}{T}\times 100$
where $t_{1},t_{2},\ldots,t_{n}$ represent the segments with missing data, and $T$ is the total audio duration.
For example, Figure 1 illustrates an instance with a 75% MDP, where up to two channels are missing at any given time. In time segments $t_{1}$ and $t_{2}$, which account for 75% of the total duration, one or more channels are missing data. In contrast, segment $t_{3}$ has no missing data across any channels. As a result, the MDP in this example is 75%.
Figure 1. Example of raw audio with 75% missing data.
Figure 2. Example of the missing channels with RF harvester.
Missing Channels in Intermittent Systems. Intermittent power systems, such as those powered by RF energy harvesters, frequently experience unpredictable power fluctuations that can lead to significant data loss in multi-channel microphone arrays. These fluctuations are caused by external factors, such as uncontrolled human traffic between the RF transmitter and receiver, which dynamically affects the system’s ability to maintain consistent power. As a result, some microphones intermittently fail to collect data, leading to incomplete or corrupted information across the array.
Figure 2 illustrates energy traces from an RF-harvested system over a 120-second period, showing how fluctuations in available energy result in a 56.67% MDP. This scenario demonstrates how missing data affects nearly 60% of the total audio, with varying numbers of channels missing at any given time.
In Section 7, we evaluate the impact of stochastic power variations on sound source localization. The results highlight the need for robust techniques like LOCUS to handle missing data in real-world applications, where power instability is common, and traditional methods often fail.
Impact of Missing Channels on SSL. There is extensive research on sound source localization (SSL) with multichannel microphones, both in simulated and real-world environments (Evers et al., 2020; Tan et al., 2021). However, while some studies address noise and reverberation, the impact of missing channel data remains underexplored.
Popular SSL input features such as Mel-frequency cepstral coefficients (MFCC) and Generalized Cross-Correlation (GCC) are highly sensitive to missing channel data, distorting feature distributions and degrading DNN performance. Figure 3 shows how MFCC and GCC change with and without missing data. It is evident how missing information distorts the features and makes them less reliable for downstream DNN tasks.To quantify this effect, we evaluate an existing SSL algorithm (Shimada et al., 2021; Adavanne et al., 2018) on the two datasets (Section 5.1), with varying levels of MDP. Figure 4 shows how DoA error increases as MDP rises, emphasizing the need for robust methods to address missing data.
(a) Without missing channels.
(b) With missing channels.
Figure 3. Mel-frequency cepstrum and Generalized Cross-Correlation of clean and corrupted audio.
Figure 4. Effect of MDP on the $E_{DoA}$
Figure 5. Effect of trivial methods on the $E_{DoA}$
Challenges in Replacing Corrupted Channels. Given the high correlation between closely spaced microphones, one might assume that excluding a faulty channel is straightforward. However, experiments show that this assumption is not effective in practice.
We conducted an experiment with two scenarios with a 75% MDP. First, we replaced missing data in faulty channels with information from the most correlated channel ($REP_{corr}$). In the second scenario, we excluded faulty channels altogether ($EXC$) and retrained the DNN to work only with the remaining active channels.
Figure  5 shows that both methods increased DoA error, by 46.34% and 63.02%. These findings demonstrate that neither detection nor retraining with fewer channels is sufficient to handle the dynamic nature of intermittent power systems, where missing data varies unpredictably. Traditional approaches fail to manage this challenge effectively in real-world applications. A more comprehensive, information entropy-based approach is needed to address missing data effectively, without requiring prior knowledge of corrupted microphones or relying on time-domain recovery models.
3. Problem Formulation and Mathematical Foundation
This section formulates the problem and provides the mathematical foundation of the proposed LOCUS.
First, we revisit the definitions of Full Rank and Low Rank Matrices. A full rank matrix $F$ has all its rows and columns linearly independent. In the context of feature extraction for sound source localization, this means that all channels provide complete and accurate information, ensuring no data loss. Conversely, a low-rank matrix $\tilde{F}$ has rows or columns that are linearly dependent, indicating missing or corrupted information in one or more channels. This loss or corruption of data reduces the effectiveness of the sound source localization algorithm.
Consider $X^{n\times T}$ as the raw time-domain data representing the input multichannel audio from a microphone array. $X^{n\times T}$ is a set of $n$ random variables , $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, where $n$ is the number of input channels and $T$ is the time duration of each channel. A feature extraction function, $G$, maps the time-domain data into the feature domain, $F^{n\times D}=G(X)$, where $D$ is the length of each feature.
When all information from all channels is present, $F$ is considered a full rank matrix. However, when one or more input channels have missing or corrupted elements, $X^{n\times T}$ changes to $\tilde{X}^{n^{\prime}\times T}$, where $n^{\prime}$ is less than $n$. As a result, the feature extraction function generates $\tilde{F}^{n^{\prime}\times D}=G(\tilde{X})$. In this paper, we focus on recovering $\hat{F}$ from $\tilde{F}$, where $\hat{F}\approx F$.
To recover the missing information, we use interpolation based on the existing information in $\tilde{F}$. However, during interpolation, the existing information in the interpolated feature map, $\bar{\bar{F}}$, may get polluted, introducing additional errors. To eliminate this residual error, we use the less-affected input information as feedback to the estimated feature.
Since it is unknown which portions of $\tilde{F}$ contain available information, we first identify the information content of $\tilde{F}$. Given that $\tilde{F}$ is a function of $\tilde{X}$, the information entropy at the $(j,k)^{\text{th}}$ element of $\tilde{F}$ is formulated as:
(1) $I_{j,k}(\tilde{F})=\sigma(-\tilde{F}_{j,k}\log\tilde{F}_{j,k})$
Here, the sigmoid function $\sigma$ in Equation 1 normalizes the entropy value between 0 and 1. $I_{j,k}(\tilde{F})$ is a real number between 0 and 1, where 1 represents the presence of all information in the $(j,k)^{\text{th}}$ element of $\tilde{F}$, while 0 indicates a complete absence of information in that element. Any value between 0 and 1 implies fractional missing information. For a full-rank feature map $F$, every element of $I(F)$ is 1.
The recovered feature set, $\hat{F}$, can be written as:
(2) $\hat{F}=\bar{\bar{F}}+I(\tilde{F})\otimes(\bar{\bar{F}}-\tilde{F})$
Here, $\otimes$ represents element-wise multiplication, and $I(\tilde{F})\otimes(\bar{\bar{F}}-\tilde{F})$ represents the feedback from the available information of $\tilde{F}$. We can rewrite Equation 2 as:
(3) $\hat{F}=I(\tilde{F})\otimes\tilde{F}+(1-I(\tilde{F}))\otimes\bar{\bar{F}}$
The first term represents the information content of $\tilde{F}$ written in terms of $\tilde{F}$ and the element-wise information entropy $I(\tilde{F})$. The second term represents the interpolated missing information extracted from $\bar{\bar{F}}$. This paper aims to determine the estimated full-rank feature matrix, $\hat{F}$, from a given low-rank feature matrix $\tilde{F}$.
4. LOCUS: Entropy-Guided Feature Restoration Framework
Figure 6. Architecture of the proposed LOCUS. The highlighted components are the contributions of this paper.
Our proposed LOCUS recovers the perturbed information of the low-rank feature matrix, $\tilde{F}$ (Equation 3), and estimates the full-rank feature matrix, $\hat{F}\approx F$. Figure 6 shows the architecture of LOCUS, which sits between the feature extraction module and the DNN-based sound source localization (SSL) task. LOCUS has three main components: (1) Early Attention (InFo), which estimates the element-wise information entropy, $I(\tilde{F})$, indicating the amount of available information in $\tilde{F}$; (2) Deep Conditional Interpolator (LaFS), which interpolates the feature set, $\bar{\bar{F}}$, from the available low-rank feature matrix, $\tilde{F}$; and (3) Guided Replacement (GRep), which combines the available information from the original feature set with the interpolated feature values guided by element-wise information entropy. Equation 3 governs this combining process. The resultant estimated full-rank feature set, $\hat{F}$, is then fed to the downstream SSL task for inference.
4.1. Information-Weighted Focus (InFo)
The purpose of Information-Weighted Focus (InFo) module is to estimate the information entropy, $I(\tilde{F})$, for each element of the low-rank feature set $\tilde{F}$. The InFo module estimates the channel and spatial information entropy (Park et al., 2018) by giving higher attention to elements with more available information. The input to InFo is the low-rank feature set $\tilde{F}\in\mathbb{R}^{d_{\tilde{X}}}$, where $d_{\tilde{X}}$ is the dimension of $\tilde{F}$. The left-hand side of Figure 6 shows the proposed early attention architecture, consisting of two main components.
(a) Retrieval of available information matrix from low-rank feature $\tilde{F}$. Here, the yellow segment signifies the missing information content in $\tilde{F}$.
(b) Retrieval of available $\hat{F}_{Z|\tilde{X}}$ from interpolated feature $\bar{\bar{F}}$. Here, the dotted green sections signify the region where error is introduced in the process of interpolation.
(c) Estimating full-rank feature $\hat{F}$ with $\hat{F}_{\tilde{X}}$ and $\hat{F}_{Z|\tilde{X}}$.
Figure 7. Step-by-step workflow of Guided Replacement (GRep) to estimate the full-rank feature $\hat{F}$ from the low-rank feature $\tilde{F}$.
The first component of InFo is the channel information branch, $f_{\gamma}(\tilde{F})$, parameterized with $\gamma$. This branch performs average pooling followed by a multi-layer perceptron to extract the channel information entropy. Average pooling aggregates information across all indices, enabling the model to learn more robust features while accurately depicting the overall strength of a feature. The multi-layer perceptron consists of three fully connected layers with batch normalization and rectified linear unit (ReLU) activation after the first two layers. This branch outputs a vector $I_{c}=f_{\gamma}(\tilde{F})$ that measures global information for each channel.
The second component of InFo is the point-wise information branch, $f_{\lambda}(\tilde{F})$, parameterized with $\lambda$. The input $\tilde{F}$ first passes through a convolution layer with a kernel size of $(1\times 1)$ and a channel reduction factor of $r$. It then goes through $n$ dilated convolution layers with a kernel size of $(3\times 3)$, dilation value $d$, batch normalization, and ReLU activation. In our architecture, $n=3$ and $d=2$. The dilated convolution increases the receptive field, enabling us to leverage contextual information. Finally, this contextual information passes through another convolutional layer with a kernel size of $(1\times 1)$ and outputs $I_{s}=f_{\lambda}(\tilde{F})$, which estimates the point-wise information of the input feature.
Next, we broadcast $I_{s}$ to $\mathbb{R}^{d_{\tilde{X}}}$ to match the dimension of $\tilde{F}$ and perform element-wise addition on $I_{c}$ and $I_{s}$ to estimate the information entropy. Finally, the sigmoid function, $\sigma$, maps the results between 0 and 1, where a value closer to 1 indicates higher information availability and vice versa. Therefore, the information entropy is defined as:
(4) $I(\tilde{F})=\sigma(I_{c}+I_{s})$
4.2. Latent Feature Synthesizer (LaFS)
The Latent Feature Synthesizer (LaFS) estimates a full-rank feature set, $\bar{\bar{F}}$, from the available low-rank feature set $\tilde{F}$ by exploiting relations among the available channels. The right-hand side of Figure 6 shows the architecture of LaFS, which takes $\tilde{F}$ as input and provides the interpolated feature $\bar{\bar{F}}$ as output. We choose an auto-encoder for estimating $\bar{\bar{F}}$ as it can learn compressed intermediate representations. Auto-encoders can efficiently interpolate missing information by semantically mixing characteristics from data (Berthelot et al., 2018) and are used in latent variable generation (Dumoulin et al., 2016; Ha and Eck, 2017).
LaFS contains two paths: (1) contraction and (2) expansion. The contraction path has $5$ downsampling blocks, each consisting of a convolution layer (kernel size = $5\times 3$), ReLU activation, and batch normalization. The output of the contraction path is the latent intermediate representation $F_{con}\in\mathbb{R}^{d_{c}}$, where $d_{c}$ represents the intermediate feature dimension.
The expansion path passes the input $F_{con}$ through 5 consecutive upsampling blocks, each consisting of transposed convolution with batch normalization and leaky-ReLU activation. The final convolution uses sigmoid activation without batch normalization. After each upsampling block, we halve the number of output feature filters, setting the last layer’s output filter number to match the shape of the input feature map. The expansion path returns the interpolated feature matrix, $\bar{\bar{F}}=f_{\phi}(\tilde{F})$, where $\bar{\bar{F}}\in\mathbb{R}^{d_{z}}$.
4.3. Guided Replacement (GRep)
The Guided Replacement (GRep) module recovers the full-rank feature $\hat{F}$ using Equation 3. Figure 7a depicts this information retrieval process. First, we perform element-wise multiplication between the information entropy $I(\tilde{F})$ and the low-rank feature set $\tilde{F}$ to retrieve the available information in the low-rank feature set $\hat{F}_{\tilde{X}}$.
Next, we find the missing information matrix $\hat{F}_{Z|\tilde{X}}$ from the interpolated feature $\bar{\bar{F}}$. Here, $Z$ represents the missing data in $\tilde{S}$. As shown in Figure 7b, we take the complement of the element-wise information matrix $I(\tilde{F})$ and multiply it element-wise with $\bar{\bar{F}}$ to find the missing information, disregarding regions (dotted darker green) where errors might have been introduced during interpolation.
Finally, GRep utilizes the available information matrix $\hat{F}_{\tilde{X}}$ and the missing information matrix $\hat{F}_{Z|\tilde{X}}$ to estimate the full-rank feature $\hat{F}$ from Equation 3 (Figure 7c).
5. Experimental Setup and Implementation Details
This section details the dataset, data preparation, implementation, baselines, and evaluation metrics for assessing LOCUS. Our implementation can be found here (Hidethyself, [n. d.]).
Table 1. Dataset Description.
{tblr}
cells = c, row1 = Mercury, row3 = Mercury, Dataset & Duration $F_{s}$(Hz) No of samples Train-Val-Test
DCASE 6 Hour 24000 600 400-100-100
LargeSet 50 hour 44100 30000 21000-3000-6000
5.1. Dataset
To evaluate LOCUS, we utilized two distinct multichannel sound source localization datasets, each offering unique challenges and characteristics. These datasets are summarized in Table 1 and detailed below.
Dataset 1: DCASE. The first dataset is DCASE2021 Task 4 (Politis et al., 2021), comprising 600 four-channel, one-minute spatial recordings at 24 kHz. This polyphonic dataset includes up to four simultaneous sound sources, 12 different acoustic events, and provides Direction of Arrival (DoA) for each source. It also features directional interference and multi-channel ambient noise, with SNR levels ranging from 6 dB to 30 dB. The dataset is split into 400 development and 200 evaluation samples.
Dataset 2: LargeSet. While large single-channel acoustic datasets like Google Audioset (Gemmeke et al., 2017) with over 2 million annotated audio clips exist for sound event detection, they do not suffice for multichannel sound source localization. To address this gap, we created and published a 50-hour, six-channel audio dataset (Authors, 2023). This dataset includes 180,000 one-second recordings at 44.1 kHz across 10 different environments (Table 2). It is monophonic, ensuring no overlapping sound sources to isolate the impact of missing information.
Simulation Details. Using the PyRoomacoustics (Scheibler et al., 2018), we position a spherical microphone array (4.25 cm radius, six microphones) at the room’s geometric center to capture realistic spatial audio data. Sound source locations are randomly selected to simulate acoustic features, including room impulse responses (RIR) and noise. Different materials emulate diverse RIRs, with RT60 measuring the decay of reverberated sound. The UrbanSound8K dataset (Salamon et al., 2014) introduces real-world noise, with the signal-to-noise ratio (SNR) randomly varied to maintain realism. Further details are available in (Authors, 2023).
Figure 8. LargeSet Simulation Environment
{tblr}
colspec=X[0]X, rowodd = Mercury, Dimension & Description
$10\times 7.5\times 3.5$ unpainted concrete - carpet, open cell foam - reverb chamber
$10\times 7.5\times 3.5$ plasterboard - cocos fibre roll - ceramic tiles
$15\times 8\times 5$ wooden lining - linoleum on concrete - lime wash
$15\times 8\times 5$ hard surface - carpet rubber - brickwork
$10\times 7.5\times 3.5$ plasterboard - carpet hairy - wooden lining
$10\times 7.5\times 3.5$ lime wash - carpet, closed cell foam - plasterboard
$15\times 8\times 5$ lime wash - felt - brick wall
$15\times 8\times 5$ lime wash - carpet cotton - concrete
$7.5\times 4\times 3$ plasterboard - carpet thin - rough concrete
$7.5\times 4\times 3$ concrete - carpet tufted - hard surface
5.2. Data Preparation
Feature Extraction. In our approach to sound source localization, we utilize two key acoustic features: Mel-Frequency Cepstral Coefficients (MFCC) and Generalized Cross Correlation (GCC). We extract these features from the raw audio data using a 1024-point Fast Fourier Transform (FFT) with a 40 ms window length and a 20 ms hop length at sampling rates of 24 kHz and 44.1 kHz. MFCC, calculated from 64 Mel bands, captures magnitude-based audio information, while GCC represents the relative time difference and phase between audio channels, crucial for determining the spatial origin of sound. We chose these features because MFCC provides important spectral details for distinguishing sound sources, while GCC captures spatial timing differences, enabling accurate localization. After truncating the GCC sequences to match the length of the Mel bins, we combine the two features into a feature matrix, $\tilde{F}$, which serves as input to deep learning pipeline for robust sound source localization. Integrating MFCC and GCC features enhances the network’s ability to accurately localize sound sources in complex acoustic environments.
Missing Channel Perturbation To prepare data with missing microphone channels, we randomly select $m\leq\frac{n}{2}$ channels from a total of $n$ microphones, where $m$ represents the number of malfunctioning or missing channels. For each Missing Data Percentage (MDP), we replace $p\%$ of the original audio with Gaussian noise, characterized by a mean of 0 and a standard deviation of 1. This noise simulates signal distortion or degraded microphone output. In addition, we simulate channel failures by randomly dropping entire channels for certain time frames (100% MDP), effectively mimicking intermittent outages where all microphones temporarily stop recording.
These perturbation techniques, including the introduction of Gaussian noise and random channel drops, ensure the model can handle real-world scenarios where microphone channels may fail or capture distorted signals.
5.3. Implementation Details
Training Pipeline of LOCUS. We train LOCUS jointly with the DNN-based sound source localization (SSL) task. The loss function for the SSL task, denoted as $\mathcal{L}_{DT}$, backpropagates through the SSL network and the channel and point-wise information branches ($f_{\gamma}$ and $f_{\lambda}$) of InFo. This ensures that the SSL task benefits from the enriched features provided by these branches.
However, the LaFS branch ($f_{\phi}$) has a separate training objective which is to estimate the input features. The loss function for DCI, $\mathcal{L}_{DCI}$, is defined as the squared $L_{2}$ distance between the original feature matrix $F$ and the reconstructed feature matrix $\bar{\bar{F}}$:
(5) $\mathcal{L}_{DCI}=||F-\bar{\bar{F}}||^{2}$
This loss function ensures that the reconstructed features closely match the original features, promoting consistency and robustness in the data representation.
To optimize the training process, we use two different optimizers: one for training InFo and the SSL task with $\mathcal{L}_{DT}$, and another for training the LaFS branch with $\mathcal{L}_{DCI}$. This dual-optimizer approach allows us to fine-tune each component of the system independently, ensuring that both the SSL task and the LaFS branch achieve optimal performance.
Sound Source Localization Pipeline.
We implement two DNN-based SLL solutions to estimate the Direction of Arrival (DoA) of sound sources using multi-channel audio – SELDNet (Shimada et al., 2021; Adavanne et al., 2018) and SALSA (Nguyen et al., 2022). SELDNet and SALSA both use MFCC as described in Section 5.2.
SELDNet is a Convolutional Recurrent Neural Network (CRNN) (Cao et al., 2019), combining CNNs for spatial feature extraction and RNNs for temporal dependencies. A fully connected (FC) layer estimates source locations in Cartesian coordinates, providing precise DoA estimates.
SALSA is an advanced SSL algorithm featuring a CRNN model with a CNN backbone based on the PANN ResNet22 (Kong et al., 2020) architecture for audio tagging. It includes a two-layer Bidirectional GRU and fully connected output layers. The algorithm adapts to various input features by adjusting the number of input channels in the first convolutional layer. During inference, sound classes with probabilities above the detection threshold are considered active, and their corresponding DOAs are selected.
5.4. Evaluation Baseline
We rigorously evaluate LOCUS by benchmarking it against four established imputation methods and one deep interpolation technique. The selected baselines encompass a spectrum of strategies, from straightforward imputation techniques to sophisticated neural network models, ensuring a thorough and robust comparison. This comprehensive evaluation framework allows us to precisely assess the effectiveness and adaptability of LOCUS in handling missing data scenarios, highlighting its strengths and potential advantages over existing solutions.
Imputation Methods. We benchmark LOCUS against three popular signal processing and two state-of-the-art deep learning imputation methods, one of which is specifically designed for acoustic signals. Each method is selected for its unique approach and effectiveness in handling missing data. Additionally, we evaluate each method on both pre-trained and retrained SSL networks.
$\bullet$ Mean Imputation ($Mean$). Missing channel information is replaced with the mean of the available channels’ information (Allison, 2001). This method is chosen for its simplicity and widespread use as a baseline in missing data problems.
$\bullet$ Hot Deck Imputation ($HotDeck$). This method fills in missing channel information using data from the most correlated channels (Little and Rubin, 2019). This method leverages the inherent similarity between channels to estimate missing data, offering a more informed and context-aware alternative to mean imputation.
$\bullet$ Probabilistic Imputation ($Prob$). Missing channel information is estimated using probabilistic methods (Fekade et al., 2018). By incorporating probabilistic models, it can better handle the complexity and randomness inherent in real-world datasets, making it a robust choice for imputation.
$\bullet$ Deep Interpolation ($AutoEnc$). We leverage the interpolation ability of autoencoders (AutoEnc) to interpolate the complex missing information and then use the interpolated feature as the input for SSL network. We compare against this method to evaluate the effectiveness of deep learning-based interpolation in handling missing data, ensuring that our model’s performance is benchmarked against advanced techniques that utilize neural networks for data reconstruction.
$\bullet$ $SoundSeive$. We apply a state-of-the-art hierarchical interpolation method for acoustic signals, as proposed by SoundSeive (Monjur et al., 2023). Unlike previous time-domain imputation methods, this approach leverages deep learning techniques in the Short-Time Fourier Transform (STFT) domain. By working in the frequency domain, $SoundSeive$ effectively captures and reconstructs complex spectral patterns, leading to a more accurate and robust imputation of missing data. We validate this method to benchmark LOCUS against cutting-edge deep learning approaches, ensuring our model’s performance is compared to the best techniques in both time and frequency domains.
Evaluation Networks. We evaluate each imputation method on two types of networks – (1) pretrained and (2) retrained.
[…]
$\bullet$ Retrained ($ReTrain$). The SSL network is retrained on corrupt data ($Corrupt$) or imputed data ($mean$, $HotDeck$, SoundSieve, and $AutoEnc$) with missing channels for $p$of the time (e.g., MDP=$75$%). Retraining with corrupt data evaluates the model’s performance when it is specifically trained to handle missing data, providing a more robust comparison. It demonstrates whether training the model with data that includes missing channels is sufficient. On the other hand, retraining with imputed data Missing data is imputed data evaluates the effectiveness of these strategies in conjunction with retraining.
5.5. Evaluation Metrics
We evaluate the performance of our sound event localization on the Degree of Arrival Estimation Error ($E_{DoA}$) of the entire dataset. $E_{DoA}$ is a suitable metric for sound source localization because it directly measures the angular difference between the estimated and true locations of sound sources. This angular error provides a clear indication of how accurately the localization system can determine the direction of incoming sounds, which is crucial for applications that rely on precise spatial audio information.
6. Results
This section begins by evaluating LOCUS on various This section evaluates LOCUS on various localization algorithms, comparing it to a state-of-the-art method to show current limitations. We analyze performance against baselines under varying missing information and conduct an ablation study to assess LOCUS’s components. Finally, we test LOCUS in a different sensor domain to show its generalizability. Section 7 further evaluates LOCUS in real-world environments.
Figure 9. Comparison among different imputation methods with pre-trained models at $MDP=75\%$.
Figure 10. Comparison among different retrained methods at $MDP=75\%$.
Figure 11. Computational complexity and overhead analysis of LOCUS, SELDNet, and SALSA.
Model
Number of
Parameter
FLOPS
Execution
Time
Power
Consumption
SELDNet 595K 20B 1.04s 120.04J
LOCUS with
SELDNet
3,686K 28B 1.08s 121.64J
SALSA 20,313M 169B 1.13s 123.96J
LOCUS with
SALSA
20,560M 178B 1.19s 132.76J
6.1. Evaluation on Different Sound Source Localization Algorithms
We evaluate LOCUS on two sound source localization algorithms – SELDNet and SALSA. Figure 15 illustrates that with no missing information, SALSA and SELDNet achieve $E_{DoA}$ values of 12.74 and 44.7 degrees, respectively. However, when MDP is 75% or 75% time-frame are missing channels, the $E_{DoA}$ increases by 38.74 degrees for SALSA and 27.07 degrees for SELDNet. The introduction of LOCUS reduces these errors by 10.34 and 7.70 degrees, respectively.
Table 11 provides a complexity and overhead analysis for SALSA and SELDNet, both with and without LOCUS. It highlights that SALSA has 34,093.87×34,093.87$\times$34 , 093.87 × more parameters and requires 8.45×8.45$\times$8.45 × more floating-point operations per second (FLOPS) compared to SELDNet. Figure 15 shows that, when combined with LOCUS, the simpler SELDNet (with only $595K$ parameters) achieves an $E_{DoA}$ that is $14.01$ degrees lower than SALSA. Moreover, LOCUS introduces only a minimal runtime overhead of $3.45-5.19\%$ and a slight increase in power consumption of $7.12-7.33\%$, which remains lower than SALSA alone.
6.2. Comparison with Baseline Algorithms
We analyze the performance of LOCUS against the baseline algorithms, as detailed in Section 5.4, using both the DCASE and LargeSet datasets.
Evaluation on Pre-trained SSL Network Figure 9 shows that without any imputation the $corrupt$ data suffers from $70.70$ and $45.52$ degree $E_{DoA}$ on DCASE and LargeSet, respectively, which is $47.89-36.95\%$ more than when all information is present. Three time-domain imputation techniques, $Mean$, $HotDeck$, and $Prob$, reduces this $E_{DoA}$ to $50.14$, $61.91$, and $159.5$ degree for DCASE and $E_{DoA}$ $55.89$, $55.68$, and $87.88$ degree for LargeSet. Imputation performs worse than the $Corrupt$ baseline because it disrupts the inherent time-frequency structure of acoustic signals.
$AutoEnc$ has $66.30$ and $52.30$ degree $E_{DoA}$ for DCASE and LargeSet, respectively which is better than $corrupt$ but still $36.73-57.93\%$ more than when all information is present. The imputation targeted for Acoustic signals, $SoundSieve$ improves performance over both $corrupt$ and $AutoEnc$ by reducing the $E_{DoA}$ to $61.40$ and $43.42$ degree which is still $31.59-49.33\%$ lower than when all information is present.
On the contrary, LOCUS uses the information estimation, recovers the missing features, and reduces $E_{DoA}$ to $36.90$ and $28.62$ degree. Which is a $13.95-76.72\%$ improvement over all the baselines on DCASE. LOCUS also outperforms all the compared baselines $32.55-67.04\%$ on LargeSet.
Evaluation on Retrained SSL Network Figure 10 compares all retrained baselines with the proposed LOCUS model. Retraining the models with corrupted data improves their overall performance compared to pre-trained models. Although retrained models such as $HotDeck$, $Mean$, and $ReTrain$ reduce $E_{DoA}$, their error rates remain $8.33-29.03\%$ higher than when full information is available on DCASE and $29.26-24.86\%$ on LargeSet. Further retrained models, $AutoEnc$ and $SoundSeive$, achieve an additional reduction in $E_{DoA}$ by $2.08-9.43\%$, yet they still perform $16.98-6.38\%$ worse than fully informed DCASE and $37.18-33.33\%$ worse on LargeSet. By contrast, LOCUS leverages mutual information across multiple channels, recovering performance by $15.19\%$ on DCASE and $24.13\%$ on LargeSet, respectively.
6.3. Effect of Missing Data Percentage
Figure 12. Performance of SELDNet, $RTrain_{p}$ and LOCUS trained with $p\%$ MDP.
Figure  12 compares the $E_{DoA}$ of LOCUS, $PreTrain$, and $ReTrain$ when we have MDP $=$ $25\%,50\%,75\%,$ and $100\%$, which is corrupt data time-frames, on DCASE and LargeSet. LOCUS achieves $9.6-51.5$ and $0.64-25.4$ degrees lower $E_{DoA}$ than $PreTrain$ on DCASE and LargeSet, respectively. With a higher percentage of missing information, $PreTrain$’s performance degrades significantly, while LOCUS consistently retrieves missing information, maintaining variances of $4.39$ and $7.89$ degrees in $E_{DoA}$ for DCASE and LargeSet. Compared to the retrained network with corrupt data ($ReTrain$), LOCUS achieves $9.6-26.1$ and $-0.05-1.78$ degrees lower $E_{DoA}$ on DCASE and LargeSet, respectively. Both $PreTrain$ and $ReTrain$ show significant increases in $E_{DoA}$ with higher MDP, as spatial relationships among input streams are crucial for localization tasks. The improvement is less pronounced for LargeSet compared to DCASE because LargeSet is monophonic with minimal MDP impact, whereas the polyphonic DCASE dataset experiences more corruption, highlighting LOCUS’s performance boost.
Figure 13. Comparison of LOCUS with SALSA at MDP=75%
Figure 14. Improvement of $E_{DoA}$ from $RTrain_{p}$ on DCASE with $MDP=\{10,25,50,75,100\}\%$.
Figure 15. Result of WiFi localization at $MDP=75\%$.
6.4. Ablation Study
This section evaluates the contribution of each component of LOCUS using the DCASE dataset and SELDNet SSL network.
Effect of InFo. To understand the impact of InFo, we evaluate LOCUS without LaFS, leaving only InFo and the SSL network. Figure 15 shows that this configuration reduces the $E_{DoA}$ by $22.85\pm 10.57\%$ compared to retrained network with corrupt data. Adding InFo enhances performance by focusing attention on the elements of the feature matrix ($\tilde{F}$) that contain more information. However, without LaFS to estimate the values of the missing elements in the full-rank version, $F$, this setup performs $14.05\pm 8.92\%$ worse than the full LOCUS implementation.
Effect of LaFS. Next, we investigate the effect of LaFS alone by removing InFo from LOCUS. This examination reveals the ability of an interpolator to recover complex missing information and the extent of additional errors introduced during interpolation. Figure 15 shows that LaFS alone reduces $E_{DoA}$ by $16.22\pm 8.87\%$ compared to retrained network with corrupt data. However, while LaFS estimates the full-rank features ($\bar{\bar{F}}$), it unintentionally introduces errors in elements that are less affected by missing data. The inclusion of InFo helps mitigate this by providing information entropy, which allows the replacement of these polluted components with their original elements, thereby reducing the overall error.
Effect of GRep. GRep in LOCUS combines the information entropy, $I(\tilde{F})$ from InFo with the interpolated full-rank feature $\bar{\bar{F}}$ from LaFS. As previously discussed, InFo estimates the available information matrix, $\hat{F}$ from ${\tilde{X}}$, while LaFS interpolates to estimate the full-rank feature, $\bar{\bar{F}}$. For feature elements in $\hat{F}_{\tilde{X}}$ significantly impacted by missing channels, GRep replaces them with interpolated elements from $\bar{\bar{F}}$. Conversely, for elements less affected by missing channels, GRep preserves the original features from $\hat{F}_{\tilde{X}}$ using InFo. This method ensures a better estimation of $\hat{F}$, enabling LOCUS to outperform both InFo and LaFS alone by $16.22\pm 8.87\%$ and $22.85\pm 10.57\%$, respectively, as shown in Figure 15. These observations confirm that retraining the baseline model alone is insufficient and that using InFo or LaFS independently is inadequate. Thus, all three components of LOCUS are crucial for recovering corrupted information.
6.5. Beyond Audio
To validate LOCUS across domains, we test it on the WiFi-based user localization algorithm DLoc (Ayyalasomayajula et al., 2020) using a large-scale WiFi CSI dataset with 16 channels (4 access points, 4 antennas each) and 75% missing data (MDP=75%). The dataset covers two indoor environments spanning 2000 sq. ft. across 8 scenarios.
Localization Network. We follow the two-step image-based feature extraction process proposed by the authors. The model architecture includes a single encoder and two decoders (consistency and location). The encoder processes input heatmaps from all access points to generate a concise representation, which is then fed into the decoders.
Result. Figure 15 shows the baseline DLoc with a 90th percentile error of 2.98 meters, while retraining ($DLoc_{retrain}$) increases the error to 3.29 meters due to varied missing data scenarios. In contrast, LOCUS reduces the error to 2.13 meters, achieving a 35.25% gain over $DLoc_{retrain}$ and 25.52% over the baseline.
7. Real World Evaluation
Figure 16. Data corruption example with energy harvester.
(a) Solar
(b) Radio Frequency
Figure 17. Comparison of LOCUS with SELDNet, and $RTrain_{75}$ with solar and radio frequency (RF) harvester.
This section presents our real-world evaluation of LOCUS across three environments under two energy scenarios, where the microphone array is powered by ambient energy.
Table 2. Environment Description.
{tblr}
width = colspec = Q[38]Q[350]Q[835], rowodd = Mercury, No & Dimension ($ft$) Description
#1 $23.5\times 46\times 10$ A large room with no window and lots of furniture and electrical equipment.
#2 $18.5\times 18\times 10$ A medium size office space with multiple furniture and large glass windows.
#3 $11\times 12.5\times 10$ A small laboratory with multiple furniture and HVAC
Energy Harvesting Setup. We employ an Ethylene Tetrafluoroethylene (ETFE) solar panel (sol, 2020) with a step-up regulator (ltc, 2015) and a Powercast harvester-transmitter system (pow, 2016a, b) to harvest radio frequency (RF) and solar energy across three environments. The solar panel is placed near a window, with intermittence caused by clouds and human shadows, while the RF transmitter and receiver are positioned on opposite walls of a corridor, experiencing interruptions from uncontrolled human traffic.
Acoustic Data Collection Setup. We utilize two MAX78000 microcontrollers to capture data from a four-channel microphone array (202, 2023), each powered by independent energy harvesters, contributing to intermittent missing channels. Since each microcontroller has only one I2S interface, two are needed to handle the array’s four-channel data. A speaker is positioned at 18 locations in 10-degree increments, with varying distances between the speaker and the array across three different environments, as detailed in Table 2. Precise measurements of the speaker’s position relative to the microphones serve as ground truth.
We collect a polyphonic dataset featuring 1-3 concurrent audio classes, using 96 one-minute segments from seven audio classes in the NIGENS database (Trowitzsch et al., 2019). The environments introduce uncontrolled noises such as HVAC sounds, conversations, and other ambient background noise. A total of 96 minutes of audio is gathered, with 72 minutes used for training and 24 minutes for evaluation. Figure 17 shows example data with a solar panel as the energy harvester.
Results. Figure 17 demonstrates that LOCUS significantly outperforms both pre-trained and retrained SELDNet, reducing $E_{DoA}$ by $25.87-60.55\%$ and $29.65-59.46\%$, respectively, in solar harvester scenarios. Interestingly, retrained SELDNet shows higher $E_{DoA}$ than the pre-trained model in environments 1 and 3. This occurs because retraining is done with $MDP=75$%, whereas the real-world introduces stochastic variations in missing information. Despite being trained on the same $MDP=75$% data, LOCUS effectively recovers stochastically missing information, highlighting the robustness of the algorithm.
In the RF harvester scenario, where sharper and more frequent energy variations occur, the challenge is more dynamic compared to solar. Figure 17b shows that RF harvesting results in more missing channels, leading to higher errors for the pre-trained SELDNet across all environments. Pre-trained and retrained SELDNet experience $118-157\%$ and $33-160\%$ higher $E_{DoA}$ than LOCUS. In contrast, LOCUS shows only $3.84-22.50\%$ more error compared to the scenario with no missing channels, further demonstrating its generalizability and robustness. Notably, in contrast to the solar harvester, retrained SELDNet has lower error than the pre-trained model in environment 1, suggesting that retraining on a fixed MDP does not adapt well to lower MDP conditions, making it less suitable for real-world applications.
8. Related Work
Intermittent System. Initial research in intermittent computing emphasized maintaining data consistency and integrity under unreliable power conditions. Systems such as Mementos (Ransford et al., 2011) introduced checkpointing strategies to periodically store system states, thus mitigating data and computation loss due to power interruptions. Although checkpointing reduces data loss, it does not fully address the issue of missing data caused by power interruptions between checkpoint intervals. Recent systems, such as Greentooth (Babatunde et al., 2024), propose resilient architectures for energy-harvesting devices to better withstand power outages. Other recent works, including FreeML (Farina et al., 2024) and SONIC (Gobieski et al., 2018), explore efficient inference of deep neural network (DNN) models on intermittent systems. In addition, (Majid et al., 2020b) demonstrate audio sensing capabilities on intermittently-powered devices; however, their system only captures audio samples of short events ( 283 ms), which remains insufficient for many real-world applications requiring extended and continuous data collection.
Missing Information Retrieval. To recover missing information from sensors, statistical methods (Jerez et al., 2010) like mean imputation (Allison, 2001), hot deck (Little and Rubin, 2019), and multiple imputations (Rubin, 1996) are often used. However, these methods focus on single-channel data and fail for long missing data sequences (Hasan et al., 2021).
In the deep learning world, missing information retrieval from a single sensor has been studied by the Natural Language Processing (NLP) and Computer Vision (CV) Community. While works in the NLP domain focus on predicting missing words and syllables in texts (Katwe et al., 2020), while CV works focus on restoring old images(Zhang et al., 2017), removing obstacles(Christiansen et al., 2016), or eliminating raindrops (Peng et al., 2020). Some other works focus on learning the input distribution (Alemi et al., 2016) to denoise the noisy inputs but can not handle the missing information. The most related work to acoustic missing information recovery is SoundSieve (Monjur et al., 2023), which recovers missing segments from single-channel audio that occurred due to energy intermittence for a sound event classification task. However, this method does not consider multiple channels and takes phase information into account for imputation. Few works on speech enhancements have looked into recovering corrupted multi-channel data (Taherian et al., 2022). However, unlike source localization, speech enhancement only partially depends on hard-to-recover spatial information and thus can not be directly used for localization tasks. Though some recent works focus on recovering missing modalities in multimodal sensing (Ma et al., 2022, 2021), no work focuses on interrelated sensors of a single modality.
Sound Source Localization. Classic sound source localization (SSL) methods, e.g., MUSIC (Gupta and Kar, 2015), independent component analysis (Noohi et al., 2013), and sparse models (Yang et al., 2018), perform poorly for under-determined scenarios. Though DNN has shown promising results for SSL (Ferguson et al., 2018; Yiwere and Rhee, 2017; Hirvonen, 2015), none of the existing works focus on missing or corrupted information.
Attention Mechanism. Attention learns to focus on essential information and ignore the rest selectively (Colombini et al., 2014). For multi-channel input, attention explores the channel characteristics and estimates the channel state information (Gao et al., 2021). Though commonly used later in a network, early attention improves performance with substantial margins (Hajavi and Etemad, 2020) motivating our InFo component.
9. Discussions and Limitations
This section describes the design choices and limitations and future directions of LOCUS.
Figure 18. Effect of retraining the downstream task.
Impact of Joint Training. LOCUS requires joint training of InFo, LaFS, and the sound source localization network to effectively recover missing information. To assess the need for retraining, we tested a setup where InFo and LaFS were added to a pre-trained, frozen SELDNet and compared it to LOCUS with a jointly trained localization network. Figure 18 shows that the pre-trained network increases $E_{DoA}$ by $14.3-103.08\%$ compared to no missing channels, while LOCUS reduces $E_{DoA}$ by $1.9-16.62\%$. This is because InFo alters the feature distribution, requiring the localization network to adapt during training, making retraining essential. Future work will explore domain generalization to potentially eliminate the need for retraining and make LOCUS a plug-and-play solution for missing information recovery.
Generative Models for Information Retrieval.. Generative models like VAEs and GANs have shown promise in recovering information, but VAEs often struggle with reconstruction accuracy and diversity due to their simplified assumptions, making them inadequate for complex, discrete data. GANs, while powerful, require large datasets and are sensitive to hyper-parameters, making them difficult to train in dynamic environments. While these models could replace the auto-encoder in LaFS, we find that a simpler, easier-to-train auto-encoder performs comparably with less overhead, making it the better choice for multi-channel information retrieval in LOCUS.
LOCUS for Multi-Sensor Systems. Multi-modal systems involve data from different sources, such as video, audio, and text, which may or may not be interrelated. In contrast, multi-sensor data, like a 12-lead ECG, collects interrelated data from multiple channels. This paper demonstrates LOCUS’s effectiveness in a multi-sensor setup and suggests its potential expansion into multi-modal domains. For example, in video conferencing, LOCUS could retrieve corrupted audio data from video and text to enhance overall quality.
Microphone Geometry Dependency. LOCUS operates on top of existing sound source localization methods and, like them, relies on the microphone-array geometry through time difference of arrival (TDoA) and triangulation. As a result, it performs poorly when used with a different geometry, requiring retraining each time the geometry changes. For example, the network trained in section 6.2 cannot be applied directly in section 7. Our future work plans to explore domain generalization techniques to address this dependency.
Large Real-World Localization Dataset. In the absence of large public datasets for sound source localization, we created and released a 50-hour simulated dataset. While useful for training DNNs, simulated data may lack the diversity and complexity of real-world conditions due to inherent assumptions. Future work will focus on developing a comprehensive real-world sound source localization dataset.
Image Recovery Limitations for Acoustic Features. While deep learning methods for image pixel recovery are effective (Zhang et al., 2017; Chen et al., 2022), they may not work for frequency domain features like GCC and MFCC. Unlike images, where neighboring pixels often share similar values, these acoustic features lack such similarity, making image recovery methods unsuitable for restoring missing acoustic channel information.
10. Conclusion
In this paper, we develop an information retrieval-based sound source localization system that can hold its performance in the presence of missing or corrupted data. We leverage the interdependence among different random variables among the channels to recover missing information. Meanwhile, classical algorithms to handle missing data points are insufficient due to their limitations, e.g., the requirement of prior knowledge and performance degradation on high and complex data distributions. LOCUS overcomes these challenges by proposing InFo, responsible to quantify how much and where the corruptions are and LaFS to interpolate them. Our evaluation shows that our proposed LOCUS can achieve up to $36.91\%$ less degree of arrival error than just retraining the localization model using data with missing channels on DCASE dataset. Our real-world evaluation shows $25.87-59.46\%$ performance gain in stochastically missing information.